After leaving Moose Jaw, Hanson drove 200 kilometers at x km/hr and then decreased his speed by 20 km/ and drove 240 additional kilomethers. Write a rational expression for his total traveling time. Evaluate the expression for x = 100.
Answer by rapaljer(4551)

D=RT
T=D/R

T(x) = D/R + D/R
T(x) = 200/x + 240/(x-20)

When x= 100,
T(100) = 200/100 + 240/(100-20)
T(100) = 200/100 + 240/80
T(100) = 2 + 3 = 5 hours

R^2 at SCC